# The Holomorphic Extension Property for $k$-du Bois Singularities

### v2

@@ 6.4. A Remark on the Functorial Pullback Morphism
Corollary 6.4 is optimal for weakly $k$-du Bois singularities. If $X$ is Cohen-Macaulay, then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for $p\leq\mathrm{codim}_{X}(\Sigma)-2$. This follows since:
@@
Recall from §6.2 that the second condition holds for isolated weakly $k$-du Bois singularities whenever $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for $p\leq k$. The first condition also holds in degree $p=k+1$ by Theorem 6.3. Since $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ agrees with the sheafification of $\Omega_{X}^{p}$ in the $h$-topology [HJ14], $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive whenever these conditions for functorial pullback hold. The following corollary of Theorem 6.3 is therefore immediate.
@@
Let $X$ be a normal complex variety with singular locus $\Sigma$, and let $\pi:\widetilde{X}\to X$ be a log-resolution of singularities. Suppose $X$ is weakly $k$-du Bois for $k\geq\mathrm{codim}_{X}(\Sigma)-1$ and that $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for some $p\leq k$. Then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p+1}$ is reflexive if and only if $\pi_{*}\Omega_{E_{i}}^{p+1}=0$ for each smooth component $E_{i}$ of the log-exceptional divisor $E$ of $\pi$.
@@ 6.5. A Remark on the Reflexivity of $\Omega_{X}^{p}$
Let $X$ be a normal complex algebraic variety. We wish to consider the $k$-du Bois and $k$-rational definitions defined in [MOPW21], [JKSY21], [MP22], [FL22a], and [FL22b]. These papers consider lci singularities, in which case the $k$-du Bois properties implies $\Omega_{X}^{p}$ is reflexive for every $p$; this is particularly special for lci singularities and restricts the codimension of the singular locus. For instance, if $Y$ is a variety with lci singularities, $\Omega_{Y}^{1}$ is reflexive if and only if $Y$ is smooth in codimension 3 [Kun86].
Understanding what happens when $\Omega_{X}^{p}$ is reflexive seems to be a difficult problem. Instead, we consider what happens when $X$ is $k$-rational for large $k$.
Theorem 6.6.
If $X$ is a complex algebraic variety with gorenstein $k$-rational singularities for $k>\frac{1}{2}\dim X$, then $X$ is smooth.
@@
Let $x\in X$. For $p\leq k$, consider the stalk $\Omega_{X,x}^{p}$. Then $\mathrm{Ext}^{j}_{\mathscr{O}_{X,x}}(\Omega_{X,x}^{p},\omega_{X,x})\cong\mathscr{H}^{j}\mathbb{D}_{X}(\underline{\Omega}_{X}^{p})_{x}=0$ for $j>0$. This implies $\Omega_{X,x}^{p}$ is projective as an $\mathscr{O}_{X,x}$-module, and the same holds for $\mathscr{F}_{p}:=\Omega_{X,x}^{p}\otimes(\omega_{X}^{*})_{x}$ since $X$ is gorenstein. There is then a trace operator
$\mathscr{F}_{p}\otimes(\mathscr{F}_{p})^{*}\to\mathscr{O}_{X,x},$
and so any minimal generating sets of $\mathscr{F}_{p}$ and $(\mathscr{F}_{p})^{*}$ have the same cardinality, i.e., $\mathrm{rk}~{}\mathscr{F}_{p}=\mathrm{rk}~{}(\mathscr{F}_{p})^{*}$.
Let $n=\dim X$. By our assumption on $k$, there is a $p<\frac{1}{2}n$ such that $\Omega_{X}^{p}\cong\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ and $\Omega_{X}^{n-p}\cong\mathbb{D}_{X}(\underline{\Omega}_{X}^{p})$. Then
$\Omega_{X}^{n-p}\cong\mathscr{H}om(\Omega_{X}^{p},\omega_{X})\cong\mathscr{H}om(\Omega_{X}^{p}\otimes\omega_{X}^{*},\mathscr{O}_{X}).$
Using again the fact that $X$ is Gorenstein, the preceding argument says that $\mathrm{rk}~{}\Omega_{X,x}^{p}=\mathrm{rk}~{}\Omega_{X,x}^{n-p}$. But $\mathrm{rk}~{}\Omega_{X,x}^{r}=\binom{e_{x}}{r}$ for any $r$, where $e_{x}$ is the embedding dimension of the singularity $(X,x)$ [Gra15, §4]. Since $e_{x}\geq n$, the equality $\binom{e_{x}}{p}=\binom{e_{x}}{n-p}$ implies that $e_{x}=n$. Therefore $X$ is smooth. ∎
Corollary 6.7.
If $X$ is a normal gorenstein variety with singular locus $\Sigma$. If $X$ is $k$-rational with $k>\frac{1}{2}\mathrm{codim}_{X}(\Sigma)$, then $X$ is smooth.
Proof.
The case of isolated singularities is covered by Theorem 6.6. The more general fact follows since a general hyperplane section of a $k$-rational singularity is $k$-rational [SVV23, Theorem A]. ∎
We remark that Corollary 6.7 was observed for the $k$-du Bois property in the case of hypersurface singularities [MOPW21, Lemma 2.2], which is equivalent to being $(k-1)$-rational [FL22b, Corollary 1.9]. We can recover this result for isolated singularities by the Proof of Theorem 6.6, since we only require:
  • $X$ is $k$-rational for some $k\leq\frac{1}{2}\dim X$,
  • $\Omega_{X}^{n-k}$ is reflexive.
While $\Omega_{X}^{n-k}$ may not be reflexive if $X$ is only $k$-du Bois, this always holds for hypersurface singularities.
7. Hodge Theory of Proper Varieties with Weakly $k$-Rational Singularities
@@ 7.1. Mixed Hodge Theory for MMP Singularities
Let $X$ be a complex algebraic variety. In his thesis, Deligne defined the notion of a mixed Hodge structure and constructed a canonical mixed Hodge structure on the cohomology of any complex algebraic variety [Del71], [Del74]. Specfically, there is an increasing weight filtration $W^{\bullet}$ on $H^{k}(X,\mathbb{Q})$ and a decreasing filtration $F^{\bullet}$ on $H^{k}(X,\mathbb{C})$ which descends to a Hodge filtration on $\mathrm{gr}_{p}^{W}H^{k}(X,\mathbb{Q})\otimes\mathbb{C}$ for each $k$. The du Bois complex is a byproduct of this construction: namely, du Bois details in [DB81] that Deligne’s construction produces a complex $\underline{\Omega}_{X}^{\bullet}$ for any algebraic variety $X$ which generates the Hodge filtration under the spectral sequence
@@
The obvious question to ask is how does the (weakly) $k$-du Bois property affect the mixed Hodge theory on the cohomology $H^{*}(X,\mathbb{Q})$ of a projective variety. Surprisingly, the vanishing cohomology of the $\underline{\Omega}_{X}^{p}$ has little control over the weight filtration:
@@
  3. (3)
    Let $Y$ be a projective hyperkähler 4-fold manifold. Suppose there exists a birational contraction $\phi:Y\to X$ of a Lagrangian submanifold $L\cong\mathbb{P}^{2}\subset Y$ to a point. Then $X$ is not weakly $k$-du Bois for $k>1$, but $H^{k}(X,\mathbb{Q})$ carries a pure Hodge structure for each $k$.
@@
More generally, we can consider the weight filtrations for the mixed Hodge structures on $H^{k}(X,\mathbb{Q})$ and $H^{2n-k}(X,\mathbb{Q}(-n))^{*}$, respectively. Since $X$ is proper, the weight filtration $W^{\bullet}$ truncates to $H^{k}(X,\mathbb{Q})$ for each $k$. The weight filtration on the dual $H^{2n-k}(X,\mathbb{Q}(-n))^{*}$ therefore is supported in higher weights (compare this to the cohomology $H^{k}(X_{\mathrm{reg}},\mathbb{Q})$), and these groups are related to the Grothendieck duals $\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ of the du Bois complex. It therefore seems better to consider how the $k$-rational property affects the weight filtration.
7.2. Purity for $k$-Rational Singularities
@@
If $X$ is a projective variety of dimension $n$ with weakly $k$-rational singularities, then the canonical mixed Hodge structure on $H^{k}(X,\mathbb{Q})$ is pure of weight $k$.
@@
generates the Hodge filtration. Note that the weight filtration $W_{\mathbb{C}}^{\bullet}$ is supported in lower weights, since $X$ is proper.
On the other hand, $\mathbb{H}^{q}(X,\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p}))\cong\mathrm{Hom}_{\mathscr{O}_{\mathrm{pt}}}(\mathbb{H}^{n-q}(X,\underline{\Omega}_{X}^{n-p}),\mathscr{O}_{\mathrm{pt}})$ by duality. Therefore $E_{1}^{p,q}$ generates the Hodge filtration on the dual mixed Hodge structure $H^{n-k}(X,\mathbb{C})^{*}$ for $k=p+q$. The weight filtration on this mixed Hodge structure is supported in higher weights. Therefore $H^{k}(X,\mathbb{Q})$ must be a pure Hodge structure. ∎
@@
If $X$ is a projective variety of dimension $n$ with weakly $k$-rational singularities, then there is a non-canonical decomposition
$H^{k}(X,\mathbb{C})=\bigoplus_{p+q=k}H^{q}(X,\pi_{*}\Omega_{\tilde{X}}^{p})$
@@
We remark that Theorem 7.2 implies something stronger than the cohomology $H^{k}(X,\mathbb{Q})$ carrying a pure Hodge structure. For instance if $k<\mathrm{codim}_{X}(\Sigma)$, where $\Sigma$ is the singular locus, the above proof shows that $H^{k}(X,\mathbb{Q})\to H^{k}(X_{\mathrm{reg}},\mathbb{Q})$ is an isomorphism. If $X$ has rational singularities, this morphism is always injective for $k=2$. If $X$ is a 3-fold with isolated singularities, then $H^{2}(X,\mathbb{Q})\to H^{2}(X_{\mathrm{reg}},\mathbb{Q})$ is an isomorphism if and only if Poincaré duality holds: therefore the defect $\sigma(X)$ of $X$ must be 0 [Kaw88, p. 97], and $X$ is $\mathbb{Q}$-factorial. This extends to higher dimensions in special cases, see for example [Tig22, Proposition 2.17].

### v3

@@ 6.4. A Remark on the Functorial Pullback Morphism
Corollary 6.4 is optimal for weakly $m$-Du Bois singularities. If $X$ is Cohen-Macaulay, then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for $p\leq\mathrm{codim}_{X}(\Sigma)-2$. This follows since:
@@
Recall from §6.2 that the second condition holds for isolated weakly $m$-Du Bois singularities whenever $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for $p\leq k$. The first condition also holds in degree $p=k+1$ by Theorem 6.3. Since $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ agrees with the sheafification of $\Omega_{X}^{p}$ in the $h$-topology [HJ14], $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive whenever these conditions for functorial pullback hold. The following corollary of Theorem 6.3 is therefore immediate.
@@
Let $X$ be a normal complex variety with singular locus $\Sigma$, and let $\pi:\widetilde{X}\to X$ be a log-resolution of singularities. Suppose $X$ is weakly $m$-Du Bois for $k\geq\mathrm{codim}_{X}(\Sigma)-1$ and that $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is reflexive for some $p\leq k$. Then $\mathscr{H}^{0}\underline{\Omega}_{X}^{p+1}$ is reflexive if and only if the fibers $H^{0}(E_{t},\Omega_{E_{t}}^{p+1}/\mathrm{tor})=0$.
@@ 6.5. A Remark on the Reflexivity of $\Omega_{X}^{p}$
Let $X$ be a normal complex algebraic variety. We wish to consider the $m$-Du Bois and $m$-rational definitions defined in [MOPW21], [JKSY21], [MP22], [FL22a], and [FL22b]. These papers consider lci singularities, in which case the $m$-Du Bois properties implies $\Omega_{X}^{p}$ is reflexive for every $p$; this is particularly special for lci singularities and restricts the codimension of the singular locus. For instance, if $Y$ is a variety with lci singularities, $\Omega_{Y}^{1}$ is reflexive if and only if $Y$ is smooth in codimension 3 [Kun86].
Understanding what happens when $\Omega_{X}^{p}$ is reflexive seems to be a difficult problem and rather restrictive. One reason for this is that the minimal generating sets of the sheaves $\Omega_{X}^{k}$ are related to the embedding dimension of $X$; in particular, a minimal generating set of the $\mathscr{O}_{X,x}$-module $\Omega_{X,x}^{k}$ has $\binom{e_{x}}{k}$ generators, where $e_{x}$ is the embedding dimension of the singularity $(X,x)$ [Gra15, §4]. In many cases, the reflexivity of $\Omega_{X}^{k}$ necessarily restricts the embedding dimension. The reflexivity of the sheaf of Kähler differentials seems to be particularly special: the only known example of a singular variety with $\Omega_{X}^{p}$ reflexive for some $p\geq 1$ are locally complete intersections. Even in this case, the higher Kähler $p$-forms will contain torsion and cotorsion [Gra15, Theorem 1.11]. Beyond this, little is known about the reflexivity of the sheaves of Kähler $p$-forms in general. We give one new result in this direction:
Proposition 6.6.
If $X$ is an $n$-fold Gorenstein variety with at worst quotient singularities, then $\Omega_{X}^{p}$ is reflexive for some $1\leq p\leq n$ if and only if $X$ is smooth.
@@
We use a description of the sheaves $\pi_{*}\Omega_{\widetilde{X}}^{p}$ used in [KS21, §10]. Let $X\subset Y$ be an embedding of $X$ into a smooth complex manifold $X$ of codimension $c$. Let $\mathcal{IC}_{X}$ be the $\mathscr{D}_{Y}$-module associated to the intersection Hodge module. The graded components of the de Rham complex with respect to the Hodge filtration are of the form
$\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})=[\Omega_{Y}^{p+c}\otimes F_{c}\mathcal{IC}_{X}\xrightarrow{\nabla}\Omega_{Y}^{p+1+c}\otimes\mathrm{gr}_{c+1}^{F}\mathcal{IC}_{X}\xrightarrow{\nabla}...\xrightarrow{\nabla}\Omega_{Y}^{n+c}\otimes\mathrm{gr}_{n-p+c}^{F}\mathcal{IC}_{X}]$
shifted by degree $-(n-p)$. Therefore, we have
$\Omega_{X}^{[p]}\cong\mathrm{ker}(\Omega_{Y}^{p+c}\otimes F_{c}\mathcal{IC}_{X}\xrightarrow{\nabla}\Omega_{Y}^{p+1+c}\otimes\mathrm{gr}_{c+1}^{F}\mathcal{IC}_{X})$
whenever $X$ has rational singularities by (3.7).
Now we use the assumption $X$ has at worst quotient singularities. Since $X$ is a rational homology manifold, the natural morphisms
$\underline{\Omega}_{X}^{p}\to\mathrm{gr}_{-p}^{F}\mathrm{DR}(\mathcal{IC}_{X})[p-n]\to\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$
are quasi-isomorphisms for every $p$; in fact, there is an isomorphism
$\mathbb{Q}_{X}[n]\xrightarrow{\sim}IC_{X}$
of perverse sheaves, and $\mathrm{gr}_{k}^{F}\mathcal{IC}_{X}=0$ for $k\geq c+1$. The above description of $\mathrm{DR}(\mathcal{IC}_{X})$ implies
(6.2) $\Omega_{X}^{[p]}\cong\Omega_{Y}^{p+c}\otimes F_{c}\mathcal{IC}_{X}.$
If we assume $\Omega_{X}^{p}$ is reflexive, then
$\Omega_{X}^{p}\cong\Omega_{Y}^{p+c}\otimes F_{c}\mathcal{IC}_{X}.$
By Claim 10.2 and the identification (6.2), this isomorphism is closed under wedging with Kähler forms on $Y$. Specifically, wedging $\Omega_{Y}^{p+c}\otimes F_{c}\mathcal{IC}_{X}$ with Kähler $(n-p)$ forms on $Y$ lands in $\Omega_{Y}^{n+c}\otimes F_{c}\mathcal{IC}_{X}\cong\omega_{X}$. By restricting to $X$, we see that
$\Omega_{X}^{n}=\Omega_{X}^{p}\wedge\Omega_{X}^{n-p}\subset\omega_{X}.$
This implies $\Omega_{X}^{n}$ is torsion-free. On the other hand, since we assume $\omega_{X}$ is a line bundle, $\Omega_{X}^{n}\to\omega_{X}$ is surjective. Since $\omega_{X}=(\Omega_{X}^{n})^{**}$, this implies $\Omega_{X}^{n}\xrightarrow{\sim}\omega_{X}$, and so $\Omega_{X}^{n}$ is locally free. This means $X$ is smooth. ∎
7. Hodge Theory of Proper Varieties with Weakly $m$-rational Singularities
@@ 7.1. Mixed Hodge Theory for MMP Singularities
Let $X$ be a complex algebraic variety. In his thesis, Deligne defined the notion of a mixed Hodge structure and constructed a canonical mixed Hodge structure on the cohomology of any complex algebraic variety [Del71], [Del74]. Specfically, there is an increasing weight filtration $W_{\bullet}$ on $H^{k}(X,\mathbb{Q})$ and a decreasing filtration $F^{\bullet}$ on $H^{k}(X,\mathbb{C})$ which descends to a Hodge filtration on $\mathrm{gr}_{p}^{W}H^{k}(X,\mathbb{Q})\otimes\mathbb{C}$ for each $k$. The Du Bois complex is a byproduct of this construction: namely, Du Bois details in [DB81] that Deligne’s construction produces a complex $\underline{\Omega}_{X}^{\bullet}$ for any algebraic variety $X$ which generates the Hodge filtration under the spectral sequence
@@
The obvious question to ask is how does the (weakly) $m$-Du Bois property affect the mixed Hodge theory on the cohomology $H^{*}(X,\mathbb{Q})$ of a projective variety. Surprisingly, the vanishing cohomology of the $\underline{\Omega}_{X}^{p}$ has little control over the weight filtration:
@@
  3. 7.3.
    Let $Y$ be a projective hyperkähler 4-fold manifold. Suppose there exists a birational contraction $\phi:Y\to X$ of a Lagrangian submanifold $L\cong\mathbb{P}^{2}\subset Y$ to a point. Then $X$ is not weakly $1$-rational [Tig24, Proposition 1.4], but $H^{k}(X,\mathbb{Q})$ carries a pure Hodge structure for each $k$.
@@
More generally, we can consider the weight filtrations for the mixed Hodge structures on $H^{k}(X,\mathbb{Q})$ and $H^{2n-k}(X,\mathbb{Q}(-n))^{*}$, respectively. Since $X$ is proper, the weight filtration $W_{\bullet}$ truncates to $H^{k}(X,\mathbb{Q})$ for each $k$. The weight filtration on the dual $H^{2n-k}(X,\mathbb{Q}(-n))^{*}$ therefore is supported in higher weights (compare this to the cohomology $H^{k}(X_{\mathrm{reg}},\mathbb{Q})$), and these groups are related to the Grothendieck duals $\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ of the Du Bois complex. It therefore seems better to consider how the $m$-rational property affects the weight filtration.
7.2. Purity for $m$-rational Singularities
@@
If $X$ is a normal and projective variety of dimension $n$ with weakly $m$-rational singularities, then the canonical mixed Hodge structure on $H^{m}(X,\mathbb{Q})$ is pure of weight $m$.
@@
generates the Hodge filtration. Note that the weight filtration $W_{\mathbb{C}}^{\bullet}$ is supported in weight $\leq n$, since $X$ is proper.
On the other hand, $\mathbb{H}^{q}(X,\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p}))\cong\mathrm{Hom}_{\mathscr{O}_{\mathrm{pt}}}(\mathbb{H}^{n-q}(X,\underline{\Omega}_{X}^{n-p}),\mathscr{O}_{\mathrm{pt}})$ by duality. Therefore $E_{1}^{p,q}$ generates the Hodge filtration on the dual mixed Hodge structure $H^{n-m}(X,\mathbb{C})^{*}$ for $m=p+q$. The weight filtration on this mixed Hodge structure is supported in weight $\geq n$. Therefore, $H^{m}(X,\mathbb{Q})$ must be a pure Hodge structure. ∎
@@
If $X$ is a projective variety of dimension $n$ with weakly $m$-rational singularities, then there is a non-canonical decomposition
$H^{m}(X,\mathbb{C})=\bigoplus_{p+q=m}H^{q}(X,\pi_{*}\Omega_{\tilde{X}}^{p})$
@@
We remark that Theorem 7.2 implies something stronger than the cohomology $H^{m}(X,\mathbb{Q})$ carrying a pure Hodge structure. For instance if $m<\mathrm{codim}_{X}(\Sigma)$, where $\Sigma$ is the singular locus, the above proof shows that $H^{m}(X,\mathbb{Q})\to H^{m}(X_{\mathrm{reg}},\mathbb{Q})$ is an isomorphism. If $X$ has rational singularities, this morphism is always injective for $m=2$. If $X$ is a 3-fold with isolated singularities, then $H^{m}(X,\mathbb{Q})\to H^{m}(X_{\mathrm{reg}},\mathbb{Q})$ is an isomorphism if and only if Poincaré duality holds: therefore the defect $\sigma(X)$ of $X$ must be 0 [Kaw88, p. 97], and $X$ is $\mathbb{Q}$-factorial. This extends to higher dimensions in special cases, see for example [Tig22, Proposition 2.17].